OpenStudy (anonymous):

Which of the following is the slope between the two points (2, 3) and (1, 4)? (a.) -1 (b.) 1 (c.) -1/3 (d.) 1/3

OpenStudy (anonymous):

To find the slope between two point, use the formula: Slope = Rise/Run = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

y2=4 , y1=3 , x2=1 , x1=2 Plug in and solve

OpenStudy (anonymous):

so we have 2=x1 3=y1 1=x2 4=y2 no?

OpenStudy (anonymous):

ok so then (4-1=3)/ (1-2=-1) and the calculator says that it is -3. Am I correct?

OpenStudy (anonymous):

oh -1/3 ok ok ok :D